Determine whether the pair of lines is parallel, perpendicular, or neither 3x+y=7, -2×+6y=6
Answer by tommyt3rd(5050)

3x+y=7,
-2×+6y=6
or
3x+y=7, m=-3
×-3y=-3, m=1/3
These have have a product of -1 so they are perpendicular


:)